OpenStudy (anonymous):

A Super Bounce Ball is dropped from a height of 64 ft. With each bounce, the ball reaches a height that is three-fourths the height of the previous bounce. After how many bounces will the ball bounce up to a height less than 9 in.? a. 3 c. 5 b. 4 d. 6

OpenStudy (anonymous):

hold on

OpenStudy (tennistar):

So, this is basically a dividing question. so you would start with 64. Divide 1/4 by 64 and you get 48.

OpenStudy (anonymous):

Please elaborate !

OpenStudy (tennistar):

Ok, so take 64. Type 1/4 into calculator on google and divide by 64. link: https://www.google.com/#q=what+is+1%2F4+of+64

OpenStudy (anonymous):

It's 16 !

OpenStudy (tennistar):

No multiply that answer by 3. 16 is 1/4 of 64. Do you get it?

OpenStudy (tennistar):

Now*

OpenStudy (tennistar):

16 x 3 = 48

OpenStudy (anonymous):

Alright ! Got it !

OpenStudy (tennistar):

You got the idea?

OpenStudy (anonymous):

a. 3 c. 5 b. 4 d. 6

OpenStudy (anonymous):

Oh shxte !! I'll try doing it now !

OpenStudy (anonymous):

Are these really the options ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

It's 4

OpenStudy (anonymous):

64 ft=> 768 inch 1st bounce= 768 x 1/4= 192 inch 2nd bounce= 192 x 1/4= 48 inch 3rd bounce= 48 x 1/4= 12 inch 4th bounce= 12 x 1/4= 3 inch so each bounce looses 3/4 of it's height !!!! I hope my explanation was clear !